Consider the following model of house pricing In(PRICE); — B + BSQFT; + B3ВED;+ BABАTН; +е

Consider the following model of house pricing In(PRICE); — B + BSQFT; + B3ВED;+ BABАTН; +е
llustrate how you can use the Unrestricted Residual Sum of Squares (URSS) and Restricted Residual Sum of Squares (RRSS) to test that the number of bedrooms and the number of bath- rooms have no effect on house prices.
Answer #1

Unrestricted model (or Full Model)

In(price) = B1B2SQFTBBed B4Bath

Restricted model (or Reduced Model)

n(price) B+B2SQFT

SSE(F) dfr SSE(R) SSE(F) F* dfR dfr

df numerator = 2

df denominator = n-k-1 = n-4

we can compare TS with critical value to test if bed and bathroom are jointly significant or not